OpenStudy (anonymous):

Solve for x. MEDAL AWARDED 3 • 4 – 3 • 7 = 3(x – 7) A. 4 B. –3 C. –10 D. 6

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

plz help me

sammixboo (sammixboo):

First multiply, what is 3 x 4?

OpenStudy (anonymous):

12

sammixboo (sammixboo):

Correct, so we have 12 - 3 x 7 = 3(x - 7) Still multiplying, what is 3 x 7?

OpenStudy (anonymous):

21

sammixboo (sammixboo):

Correct, so we have 12 - 21 = 3(x - 7) What is 12 - 21?

OpenStudy (anonymous):

-9?

sammixboo (sammixboo):

Yes, so we have -9 = 3(x - 7) and to solve this, you must simplify the equation to have 3x - 21 = -9 So we first, add 21 to boh sides of the equation 3x - + 21 = -9 + 21 3x = -9 + 21

sammixboo (sammixboo):

What is -9 + 21

OpenStudy (anonymous):

12

OpenStudy (anonymous):

is that the answer

sammixboo (sammixboo):

Correct, so we have 3x = 12

sammixboo (sammixboo):

Now we divide 3 on both sides of the equation 3x/3 = 12/3 x = 12/3

OpenStudy (anonymous):

4 is the answer?

sammixboo (sammixboo):

Correct

OpenStudy (anonymous):

OHHHHHHHH yay thnk you

sammixboo (sammixboo):

No prob!